OpenStudy (anonymous):

Factor 12x^2+13xy+3y^2

OpenStudy (4n1m0s1ty):

12x^2 +13xy + 3y^2 = 12x^2 + 9xy + 4xy + 3y^2 = 3x(4x+3y) + y (4x+3y) =(3x+y)(4x+3y)

OpenStudy (anonymous):

(3x+y)(4x+3y)

OpenStudy (anonymous):

Thankyou guys!:)

OpenStudy (anonymous):

12x^2+13xy+3y^2 (4x+3y)(3x+y)